OpenStudy (anonymous):

A student solved this problem and said the answer was 2 1/2 cups. Rhonda picked 6 2/3 cups of blackberries. Her sister picked 4 1/6 How many more cups of blackberries did Rhonda pick than her sister? Is the student's answer reasonable? A. Yes, the answer is reasonable. B. No, the answer is not reasonable. It should be about 1/2 cup. C. No, the answer is not reasonable. It should be about 4 cups. D. No, the answer is not reasonable. It should be about 10 cups.

OpenStudy (anonymous):

@shinebrightlikeadimon

OpenStudy (anonymous):

@kl0723

OpenStudy (anonymous):

@VampirePrincess

OpenStudy (anonymous):

@kittiwitti1

OpenStudy (welshfella):

6 2/3 - 4 1/6

OpenStudy (anonymous):

5/2

OpenStudy (welshfella):

convert the 2/3 to 6th's then you can do the subtraction

OpenStudy (anonymous):

guide me

OpenStudy (kittiwitti1):

\[6\frac{2}{3}-4\frac{1}{6}\]Improper fraction both then make both sixths fractions and subtract

OpenStudy (anonymous):

i think its c

OpenStudy (anonymous):

i got 5/2

OpenStudy (welshfella):

6 - 4 = 2 now you need to calculate 4/6 - 1/6 = ?

OpenStudy (welshfella):

add the 2 to the result of the subtraction of the fractions

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

1/4

OpenStudy (welshfella):

?? 1/2 or 1/4 ?

OpenStudy (anonymous):

1/2

OpenStudy (welshfella):

right now add the 2 and you have it

OpenStudy (kittiwitti1):

Mixed to improper: \[\frac{(6\times4+3)}{3}-\frac {(4\times6+1)}{6}\]

OpenStudy (kittiwitti1):

Urgh. I forgot you could solve without improper conversion....

OpenStudy (welshfella):

2 + 1/2 = ?

OpenStudy (welshfella):

yes - in this case - its easier to just subtract the whole numbers first then the fractions after

OpenStudy (kittiwitti1):

I was raised and taught that it was not logical or risky to do that ._.; *I have weird teachers*

OpenStudy (welshfella):

yes - you can be too dogmatic here its ok to do that

OpenStudy (welshfella):

its perfectly safe as 6 > 4 and 2/3 > 1/6

OpenStudy (welshfella):

well freez what do you think the answer is?

OpenStudy (anonymous):

B

OpenStudy (anonymous):

bbbbbbbbbbbbb

OpenStudy (kittiwitti1):

no

OpenStudy (kittiwitti1):

it is not b but nice try

OpenStudy (anonymous):

c

OpenStudy (anonymous):

cc

OpenStudy (kittiwitti1):

No... what answer did you get from what @welshfella told you?

OpenStudy (welshfella):

all you have to do is calculate 2 + 1/2

OpenStudy (kittiwitti1):

Yes, and the answer is neither B nor C ; ;

OpenStudy (anonymous):

a

OpenStudy (kittiwitti1):

...yes. It's A. ; ;